3. Suppose there is exactly one packet switch (or router) between a sending host and a receiving host. The transmission rates between the sending host and the switch and between the switch and the receiving host are R1 = 1Mbps and R2 = 2Mbps, respectively. Assuming that the switch uses store-and-forward packet switching, what is the total end-to-end delay to send a packet of length 18000bits? Assume that the propagation speed is 2 * 108 meters/sec, the distance between source and the router is 20,000 km, the distance between the router and the destination is 10,000 km (Ignore queuing delay and processing delay, considering transmission delay and propagation delay).

4. Review the car-caravan analogy in Section 1.4. Assume a propagation speed of 150 km/hour. Suppose the caravan travels 300 km, beginning in front of one tollbooth, passing through a second tollbooth, and finishing just after a third tollbooth. What is the end-to-end delay?

5. We are sending a 20 Mbit test file from a source host to a destination host. All links in the path between source and destination have a transmission rate of 15 Mbps. Assume that the propagation speed is 2 * 108 meters/sec, and the distance between source and destination is 20,000 km.

a) Initially suppose there is only one link between source and destination. Also suppose that the entire file is sent as one packet. The end-to-end delay (transmission delay plus propagation delay) is?

b) Now suppose there are two links between source and destination, with one router connecting the two links. Each link is 10,000 km long. Again suppose the file is sent as one packet. Suppose there is no congestion, so that the packet is transmitted onto the second link as soon as the router receives the entire packet. The end-to-end delay (transmission delay plus propagation delay) is?

c) Now suppose there are two links between source and destination, with one router connecting the two links. Each link is 10,000 km long. Furthermore, suppose that the file is broken into 4 packets, each of 5 Mbit. Ignore headers that may be added to these packets. Also ignore router processing delays. Assuming store and forward packet switching at the router, the total end-to-end delay (transmission delay plus propagation delay) is?

6. Suppose there are four links between a source and a destination. The first link has transmission rate 100 Mbps, the second link and third has transmission rate 12 Mbps, and the last link has transmission rate at 20Mbps. assuming that the only traffic in the network comes from the source, what is the throughput for a large file transfer? Remark: For simplicity, assume 1Mbit = 1000Kbit = 1000*1000 bit
Answer #1

3)

Solution:

Given that, The transmission rate between the sending host and the switch R1 = 1Mbps.
ransmission rate between the switch and the receiving host is R2 = 2Mbps

Let us assume at the time interval t0 the sending host begins to transmit.
Then at time t1 = L/R1, the sending host completes transmission and packets are at router.
t1 = 18000/1Mbps
t1 = 18000 sweconds.
At time t2 = t1 + L/R2,
t2 = 18000 + 18000/2
t2 = 27000 sec.
And the router completes transmission.
then end‐to‐end delay is nothing but t2.
= 27000 sec